Circular No. 8470

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005M AND 2005N Two apparent supernovae have been reported from unfiltered CCD

frames: 2005M by T. Puckett and D. George (cf. IAUC 8460; 0.60-m automated supernova patrol telescope), and 2005N by J. Graham and W. Li (cf. IAUC 8460; LOSS/KAIT).

SN  2005 UT  α_{2000}  δ_{2000}  Mag.  Offset
2005M Jan. 19.37 9^{h}37^{m}32.^{s}36 +23^{o}12′02′′.7 18.8 6′′.6 W, 10′′.8 S
2005N Jan. 19.58 14 04 02.61 −14 37 48.8 18.8 39′′ E, 48′′ S

Additional approximate CCD magnitudes by the respective discoverers: SN 2005M in NGC 2930, 2002 Apr. 4 UT, [20.0; 2005 Jan. 20.22, 18.2. SN 2005N in NGC 5420, 2004 July 9.21, [20.0; 2005 Jan. 20.55, 18.3.

SUPERNOVAE 2005J AND 2005K M. Modjaz, R. Kirshner, and P. Challis, Harvard-Smithsonian Center

for Astrophysics, report that a spectrum (range 350–740 nm) of SN 2005J (cf. IAUC 8467), obtained by M. Calkins on Jan. 19.40 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST), shows it to be a type-II supernova, probably within a few weeks past explosion. The spectrum consists of a blue continuum and Balmer-line P-Cyg features. Adopting the NED recession velocity of 4182 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is ∼ 10000 km/s. A noisy spectrogram of SN 2005K (IAUC 8468) taken by Calkins on Jan. 19.42 reveals it to be a type-II supernova. The spectrum shows P-Cyg features of Hα and Hβ, and the expansion velocity derived from the minimum of the Hβ line is ∼ 5000 km/s (adopting the NED recession velocity of 8151 km/s for the host galaxy).

COMET C/2005 A1 (LINEAR)

Improved parabolic orbital elements from MPEC 2005-B03.

T = 2005 Apr. 10.2414 TT  ω = 271.^{o}8882
Ω = 355.8552 }2000.0
q = 0.906662 AU  i = 74.8859

Total visual magnitude estimate by J. J. Gonzalez, Asturias, Spain (0.20-m reflector): Jan. 17.25 UT, 11.6 (coma diameter 2′.5).

2005 January 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green